Consider two normalized eigen functions and , corresponding to the same eigen value. If



Integral a1a2*dT = d



where is real



Find a normalized linear combinations of and that are orthogonal to



(a) a1



(b) a1+a2



Note: The coefficients of the linear combinations need not be real

1
Expert's answer
2022-03-23T04:13:02-0400

"\int" a1 * a2 * dT = d;

d/dT = "\int" a1 * a2

d/dT = -a2;

d/dT = "\int" (a1 + a2) * -(a2)